OpenStudy (anonymous):

Dark Trig Question?? Pleaseee??? ✏ ✐ How does changing the function from f(x) = 2 sin 4x to g(x) = 2 sin 4x + 3 affect the range of the function?

OpenStudy (anonymous):

The function shifts up 3 units, so the range changes from −2 to 2 in f(x) to 1 to 5 in g(x). The function shifts up 3 units, so the range changes from −1 to 1 in f(x) to 2 to 4 in g(x). The function shifts up 4 units, so the range changes from −2 to 2 in f(x) to 2 to 6 in g(x). The function shifts up 4 units, so the range changes from −1 to 1 in f(x) to 3 to 5 in g(x). ^Possible Answers

OpenStudy (anonymous):

@rational @dmndlife24 Do you know how to do this?

OpenStudy (anonymous):

D:

OpenStudy (anonymous):

@Er.Mohd.AMIR Do you think you can help me on this one?

OpenStudy (anonymous):

@FibonacciChick666

OpenStudy (anonymous):

@robtobey @geerky42 @TheSmartOne @kainui

OpenStudy (anonymous):

@Jaynator495 @Nnesha

OpenStudy (anonymous):

Nobody has come to my rescue... *cries*

OpenStudy (anonymous):

I think it is the first one, so I guess I will put that then. ◕╭╮◕

OpenStudy (unklerhaukus):

imagine you graph y = f(x), and also y = g(x) on the same plot. The y-values of g(x) will always be 3 more than (above) the y-values of f(x), ie g(x) = f(x) + 3,

OpenStudy (er.mohd.amir):

range of sin x is between -1 to 1 and 2 sin x is become -2 to 1 and 2 sin 4x also same and adding 3 range become -2+3 to 2+3 so 1 to 5 is ans